OpenStudy (anonymous):

3x - 5x + 8x + x = 42 I keep getting 11x but it can't divide into 42

OpenStudy (texaschic101):

3x - 5x + 8x + x = 42 12x - 5x = 42 7x = 42 x = 42/7 x = 6

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

Yes a few

OpenStudy (texaschic101):

ask away

OpenStudy (anonymous):

I thought I couldn't Jump around like you did?Thats why I didn't add the 3x I thought you couldn't change the addition signs and subtraction signs on these type of problems?

OpenStudy (texaschic101):

I didn't change the signs.....the 3x , the 8x, and the (1)x are positive and they equal 12x. Then there is the -5x....so you subtract 5x from 12x and that leaves you with 7x. You then divide by 7 and that leaves x equal to 6. Do you understand ?

OpenStudy (anonymous):

Yes I just thought I couldn't move around like you did solving the problem is what im trying to say

OpenStudy (texaschic101):

First I add all the positives together, and then I add all the negatives together, and then I subtract, or add (whatever the signs lead me to do) You can move around.

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

no problem :) Have a nice day